OpenStudy (anonymous):

During the year, the Cheung’s real estate bill included $224 for county services. Of this amount, $101 went to the highway department. What percent did the county highway department receive? Round your answer to two decimal places.

OpenStudy (anonymous):

You want to take the amount that went to the highway and divide by the amount that went to county services... that will give you what percentage went into highway... 101/224...then plug into a calculator to get the decimal

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The same way if you had $10 and put $5 in the bank. What percentage did you put in the bank. 5/10= 1/2 = 50% does that make sense?

OpenStudy (anonymous):

yeah yeah thank you

OpenStudy (anonymous):

welcome